OpenStudy (anonymous):

How do you DASH things on Differentiation eg. My tutor said this: y= Root (2x + 1) y`= (2x+1)` --------- (fraction) 2 root(2x+1) = 2 ------ 2 root ( 2x+1) HOW HOW HOW! sorry stress

OpenStudy (anonymous):

I really need help. it due tomorrow. a whole excerise

OpenStudy (anonymous):

HELPLPLP

OpenStudy (unklerhaukus):

The chan rule\[y(x)=f\big(g(x)\big)\] \[y'(x)=g'(x)f'\big(g(x)\big)\]

OpenStudy (anonymous):

No, I DON'T understand the dashing in general. Even a simple equation. I dont know understand how the numbers appear

OpenStudy (anonymous):

how does root ( 2x+1) become 2/ (2)(root of 2x+1)

OpenStudy (anonymous):

HOW?

OpenStudy (unklerhaukus):

so \[y(x)= \sqrt{2x + 1}=(2x+1)^{1/2}\] \[f(x)=\sqrt x\]\[g(x)=2x+1\] can you tell me \(f'(x)\) ?, \(g'(x)\) ?

OpenStudy (anonymous):

f` = 1x f^-1 (^ is power of) CORRECT? WRONG?

sam (.sam.):